OpenStudy (anonymous):

prove that if p<= n, then p does not divide n!+1 Please help

OpenStudy (perl):

this is the proof to the prime numbers being infinite

OpenStudy (anonymous):

Is that your question ,

terenzreignz (terenzreignz):

more like... \[\huge p < n \ \ \rightarrow \ \ \frac{n! + 1}{p}\notin \mathbb{Z}\]

terenzreignz (terenzreignz):

Okay... no need to get so edgy :D\[\huge p \le n \ \ \rightarrow \ \ \frac{n! + 1}{p}\notin \mathbb{Z}\]

terenzreignz (terenzreignz):

p = 1 instant contradiction http://tinyurl.com/3n3sxpy haha

OpenStudy (perl):

1 is not a prime

terenzreignz (terenzreignz):

oh... lol I didn't know p had to be a prime :) I just took it for a variable. My bad :D

terenzreignz (terenzreignz):

Why does p even have to be prime? Doesn't it apply for all integers greater than 1?

OpenStudy (anonymous):

yes i understand the question, but i have no idea how to start this question.. D:

terenzreignz (terenzreignz):

Familiar with modular math, @neoc? :)

OpenStudy (anonymous):

@terenzreignz yes iam familiar with modular math

terenzreignz (terenzreignz):

That is awesome :) Then you'd know that if 1 < p < n then \[\huge p\lvert n!\] right? Because \[\huge n! = n\cdot(n-1)\cdot...p ...3\cdot2\cdot1\]or something

OpenStudy (anonymous):

oh yes! thank you so much! it all make sense now :D

terenzreignz (terenzreignz):

Wow.. that escalated quickly :D No problem :)